# GaussianFocus: Constrained Attention Focus for 3D Gaussian Splatting

### v1

@@ 3.5 Training Losses
where $\nabla_{x}$ and $\nabla_{y}$ represent the gradient operator calculated using the Sobel filter, capturing edge information along the $x$ and $y$ directions. The $p_{i}$ and $\hat{p}_{i}$ represent the pixels of the ground truth image $G_{i}$ and the corresponding pixel in the rendered image $P_{i}$, Moreover, we introduce the frequency loss term to address the challenge of high-frequency detail loss. It approximates the frequency domain loss by employing gradient loss computations in the $x$ and $y$ directions for both the input and target images. This term is essential for preserving high-frequency details and is computed as:
@@
$L_{\text{Total}}=\begin{cases}(1-\lambda)L_{1}(p_{i},\hat{p}_{i})+\lambda L_{\text{D-SSIM}}(p_{i},\hat{p}_{i})+\beta L_{\text{Edge}}(p_{i},\hat{p}_{i}^{{}^{\prime}})+\eta L_{\text{Frequency}}(p_{i},\hat{p}_{i}^{{}^{\prime}}),\quad\text{every 50 iterations,}\\ (1-\lambda)L_{1}(p_{i},\hat{p}_{i})+\lambda L_{\text{D-SSIM}}(p_{i},\hat{p}_{i})+\beta L_{\text{Edge}}(p_{i},\hat{p}_{i})+\eta L_{\text{Frequency}}(p_{i},\hat{p}_{i}),\quad\text{otherwise,}\end{cases}$ (14)
where $\hat{p}_{i}^{{}^{\prime}}$ denotes the pixel in the attention-enhanced image. The $\lambda$, $\beta$ and $\eta$ are the respective weights assigned to the loss components and they are set to 0.2.
4 Experiments
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Original Res. 1/2 Res. 1/4 Res. 1/8 Res. Avg. Original Res. 1/2 Res. 1/4 Res. 1/8 Res. Avg. Original Res. 1/2 Res. 1/4 Res. 1/8 Res. Avg. NeRF 0.933 0.966 0.970 0.948 0.954 31.27 31.98 29.98 26.52 29.94 0.059 0.040 0.049 0.059 0.052 Mip-NeRF 0.960 0.968 0.970 0.960 0.965 32.50 33.00 31.20 28.10 31.20 0.044 0.030 0.035 0.051 0.040 Instant-NGP 0.963 0.968 0.965 0.946 0.961 33.05 33.10 29.80 26.45 30.60 0.046 0.036 0.048 0.072 0.051 TensoRF 0.958 0.970 0.960 0.950 0.960 32.60 32.75 30.20 26.30 30.46 0.046 0.035 0.047 0.070 0.050 Tri-MipRF 0.961 0.969 0.953 0.908 0.948 32.75 33.00 29.70 24.10 29.89 0.048 0.038 0.048 0.072 0.051 3DGS 0.973 0.952 0.868 0.761 0.889 33.50 27.10 21.60 17.80 25.00 0.032 0.022 0.068 0.118 0.060 3DGS + EWA 0.967 0.964 0.955 0.943 0.960 33.60 31.80 27.95 24.75 29.53 0.035 0.026 0.036 0.049 0.037 Mip-Splatting 0.966 0.971 0.965 0.973 0.968 33.11 33.88 31.40 27.77 31.54 0.038 0.025 0.029 0.031 0.031 Ours 0.988 0.981 0.982 0.985 0.984 33.79 33.96 31.74 28.65 32.04 0.031 0.017 0.019 0.021 0.022
Table 1: Quantitative Comparison with Baselines on the Blender Dataset [4]. All models are evaluated at four progressively lower resolutions and trained using images at original resolutions. Our method outperforms other models at 1/2, 1/4, and 1/8 resolutions and achieves comparative results at the original resolution.
4.1 Baselines
We selected Mip-Splatting [7] and 3D-GS [5] as our primary baseline due to their established state-of-the-art performance in novel view synthesis. In our evaluation, we included several other prominent techniques, such as Mip-NeRF360 [16], Mip-NeRF [34], Instant-NGP [13], Zip-NeRF [35], Scaffold-GS [8], TensoRF [10], and Tri-MipRF [51]. We also considered NeRF [4] and 3DGS + EWA [52] for further comparison. They are the most representative models.
@@ 4.1 Baselines
4.2 Datasets and Metrics
We carried out an extensive evaluation of multiple scenes sourced from publicly available datasets, including a dataset that features a division of a large scene. Specifically, we assessed our method using seven scenes drawn from Mip-NeRF360 [16], the synthetic Blender dataset [4], a Villa scene, Mill-19 dataset [46] and two scenes from Tanks&Temples [53]. The Villa scene is a self-processed dataset. The Villa scene is a self-collected dataset. The evaluation metrics we report include Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [54], and Learned Perceptual Image Patch Similarity (LPIPS) [55]. For Mip-NeRF360 and Blender datasets, we present the average values of these metrics across all scenes to provide a comprehensive overview of our approach’s performance.
4.3 Implementation
Our approach is developed on the foundation of the open-source 3DGS code [5]. Adhering to the protocol established in [5], we train our models and baselines for 30,000 iterations over all scenes, utilizing the same Gaussian density control strategy, training pipelines and hyperparameters. Furthermore, patch attention is utilized to enhance reconstruction quality every 50 iterations. We also inspect and constrain the scale matrix $S$ of Gaussian spheres every 1,000 iterations, up to the first 10,000 iterations. We set the kernel size as 0.05 and the loss weight $\lambda=0.2$.
@@ 4.4 Result Analysis
Comparison on the Blender Dataset    Following prior work [7], we trained our model on scenes at their original resolution and rendered them at four different resolutions: original, 1/2, 1/4, and 1/8. The quantitative results are detailed in Table. 1 which shows our method outperforms baselines. Our analysis includes NeRF-based [4] and 3DGS-based [5] methods that highlight consistent performance gains across all resolutions, especially at lower resolutions.
Comparison on the Mip-NeRF 360 Dataset    In our experiments, we trained models on data downsampled by a factor of eight, and then rendered images at different resolutions (1/8, 1/4, 1/2, and original resolution). As illustrated in Table. 2, our method matches prior work at the training resolution (1/8) and significantly outperforms existing state-of-the-art methods at higher resolutions (1/4, 1/2, and original). Fig. 4 demonstrates that our approach renders high-fidelity images without introducing high-frequency artifacts. This is in stark contrast to Mip-NeRF 360 [16] and Zip-NeRF [35], which tend to falter at higher resolutions due to their MLP architectures’ limitations in managing unrepresented frequencies during training. Moreover, the 3DGS method [5] often yields significant degradation artifacts due to its reliance on dilation processes. Although the 3DGS + EWA method [52] mitigates some issues, it still produces noticeable high-frequency artifacts. Our method avoids these issues and more accurately represents the ground truth. Additionally, our method effectively reduces blurred artifacts in Mip-splatting [7].
@@
Comparison on the Villa Dataset    In the Villa Dataset experiment, we compared our model’s training progression against Mip-Splatting [7], with both models trained at original resolution. Results in Fig. 5 show the performance of both models across 100, 900, 2k, and 5k iterations. Our model demonstrated significant improvements by the 900th iteration, whereas scenes from Mip-Splatting [7] remained blurry and of lower quality. Even after 5k iterations, Mip-Splatting failed to achieve the detail level of our model at 900 iterations, especially in fine features such as roofs, windows, and walls.
Figure 7: Ablation of Gaussian Patch Attention Strategy. We present an ablation study of our model trained on the Garden scene [16], comparing results at 30k iterations with and without the application of the Gaussian Patch Attention Enhancement Strategy.
Evaluation on the Large Scene Dataset    In our study, we addressed the challenges of reconstructing large scenes like small towns or city-scale environments, which are unmanageable for traditional 3DGS-based [5] and NeRF-based [4] models due to memory constraints and long optimization times. We used the Mill-19 Rubble scene [46], which had excessively noisy point clouds requiring reprocessing and selective image filtering. We subdivided the scene, which contained over 1,700 images, into 64 blocks. Each block was independently trained with 200 to 500 images. This reduced memory demands and allowed efficient parallel training in just 20 minutes. Our reconstruction results depicted in Fig. 6, show the seamless reassembly of all blocks which preserves the continuity of the large-scale scene. This method contrasts with previous models, which failed to directly reconstruct large scenes and compromised on reconstruction quality by randomly selecting a subset of images for training.
Figure 8: Ablation of Gaussian Sphere Constraints Strategy. We present an ablation study of our model trained on the Villa scene, comparing results with and without the application of the Gaussian Sphere Constraints Strategy. This strategy reduces the “air walls” problem.
@@ 4.5.1 Patch Attention Enhancement
Omitting the $L_{Frequency}$ and $L_{Edge}$ from the Patch Attention strategy leads to a significant reduction in rendering quality. As shown in Fig. 7, removing this strategy leads to noticeable degradation in rendering quality, especially in image details. Without Patch Attention, images exhibit blur effects due to high-frequency dilation issues. To quantitatively evaluate this impact, we referred to Table. 3, which compares performance metrics with and without this enhancement. The results clearly indicate improvements across all metrics when the Patch Attention strategy is employed, significantly enhancing the model’s ability to produce detailed and sharp renderings by focusing on edge information.
@@ 4.5.2 Gaussian Sphere Constraints
We assessed the importance of Gaussian Sphere Constraints by removing them from our model. As shown in Fig. 8, models rendered without these constraints exhibit oversized Gaussian spheres, which result in information loss and reduce the overall quality of the renderings. In 3D scenes, these oversized spheres often create “air walls” in detail-heavy areas. Implementing Gaussian Sphere Constraints allows us to effectively control the growth and size of these spheres, enhancing detailed depiction within the scene. The comparative images in Fig. 8, especially in the lower two layers, clearly demonstrate the loss of detail in models rendered without this strategy. These images highlight how the constrained Gaussian spheres maintain finer details, leading to more precise and realistic renderings. Additionally, as indicated in Table. 3, the inclusion of Gaussian Sphere Constraints significantly improves performance metrics.
@@ 4.5.3 Subdivision-Based Reconstruction of Large Scenes
In this work, we are unable to conduct an ablation study for our ”Subdivision-Based Reconstruction of Large Scenes” approach. This is primarily due to the limitations of existing 3DGS-based methods [5, 7, 8, 9, 49], which cannot be trained on our large-scale scene datasets without encountering out-of-memory errors. Our subdivision method is specifically engineered to facilitate the training of 3DGS-based methods on exceptionally large scenes—a feat unattainable with current state-of-the-art 3DGS-based techniques without supplementary modifications. The indispensable nature of our method implies that any modification or removal of this subdivision strategy for ablation purposes would render the application of 3DGS on large scenes entirely infeasible. This is because our approach is not simply an incremental enhancement but a crucial component that establishes the possibility of training in such extensive environments.
Figure 9: Limitation: The result of the recombined scene will contain the boundary artifacts of the result of the previous small block.
@@ 4.6 Limitation
Our model integrates the Patch Attention Enhancement feature, which substantially improves the quality of rendered images by meticulously calculating attention values. While this method enhances detail recognition and overall image fidelity, it also significantly increases the memory demands of the model. This elevated memory consumption has the potential to trigger out-of-memory errors during the training phase, particularly with complex or large-scale scenes. To address this limitation, future versions of the model could explore alternative computational methods or more efficient data structures, which might reduce the memory requirements while maintaining or even enhancing the model’s performance. Another challenge arises in the reconstruction of large scenes where the final assembly of individual blocks can lead to complications. Specifically, the boundaries of each block may overlap, causing visible disruptions in the continuity of the scene. These overlaps often manifest as clusters of disorganized Gaussian spheres at the edges, which are evident in the reconstructed images shown in Fig. 9. This not only affects the aesthetic quality of the renders but also detracts from the model’s utility in practical applications. In the future, it may be beneficial to design an algorithm that removes Gaussian spheres at the boundaries of each block. This would enhance the quality of the final assembled large scene and ensure a more natural and seamless appearance.
@@ 5 Conclusion
In this paper, we present GaussianFocus, an enhanced model derived from traditional 3D Gaussian Splatting. It features three key innovations: Patch Attention Enhancement, Gaussian Constraints Strategy and the subdivision of large-scale scenes into manageable blocks for individual training. These innovations aim to refine detail representation, enhance reconstruction quality and reduce the “air walls” problem. The approach of subdividing large scenes into manageable blocks overcomes the limitations inherent in traditional 3DGS-based methods, which struggle with extensive scenes. Experimental results demonstrate that GaussianFocus competes well with state-of-the-art methods at a single scale and excels across multiple scales, providing superior detail accuracy and reconstruction quality.

### v2

@@ 3.4 Training Losses
where $\nabla_{x}$ and $\nabla_{y}$ represent the gradient operator calculated using the Sobel filter, capturing edge information along the $x$ and $y$ directions. The variables $p_{i}$ and $\hat{p}_{i}$ represent the pixels of the ground truth image $G_{i}$ and the corresponding pixel in the rendered image $P_{i}$, respectively. Moreover, we introduce a frequency loss term to address the challenge of high-frequency detail loss. It approximates the frequency domain loss by employing gradient loss computations in the $x$ and $y$ directions for both the input and target images. This term is essential for preserving high-frequency details and is computed as:
@@
$\begin{split}L_{\text{Total}}=\begin{cases}(1-\lambda)L_{1}(p_{i},\hat{p}_{i})+\lambda L_{\text{D-SSIM}}(p_{i},\hat{p}_{i})+\\ \beta L_{\text{Edge}}(p_{i},\hat{p}_{i}^{{}^{\prime}})+\eta L_{\text{Frequency}}(p_{i},\hat{p}_{i}^{{}^{\prime}}),\quad\text{every 50 iterations,}\\ (1-\lambda)L_{1}(p_{i},\hat{p}_{i})+\lambda L_{\text{D-SSIM}}(p_{i},\hat{p}_{i})+\\ \beta L_{\text{Edge}}(p_{i},\hat{p}_{i})+\eta L_{\text{Frequency}}(p_{i},\hat{p}_{i}),\quad\text{otherwise,}\end{cases}\end{split}$ (11)
where $\hat{p}_{i}^{{}^{\prime}}$ denotes the pixel in the attention-enhanced image. The variables $\lambda$, $\beta$ and $\eta$ are the respective weights assigned to the loss components.
@@
4 Experiments
4.1 Baselines, Datasets, Metrics and Implementation
We selected Mip-Splatting [30] and 3D-GS [9] as our primary baseline due to their established state-of-the-art performance in novel view synthesis. In our evaluation, we included several other prominent techniques, such as Mip-NeRF360 [2], Instant-NGP [16], and Zip-NeRF [3]. We also considered 3DGS + EWA [32] for further comparison.
We extensively evaluated multiple scenes from publicly available datasets, including a dataset that features a division of a large scene. Specifically, we assessed our method using seven scenes drawn from Mip-NeRF360 [2], a Villa scene, Mill-19 dataset [23] and two scenes from Tanks&Temples [11]. The Villa scene is a self-processed dataset. The evaluation metrics we report include Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [26], and Learned Perceptual Image Patch Similarity (LPIPS) [31].
Following 3DGS [9], we trained our models for 30k iterations and applied patch attention every 50 iterations, and adjusted the scale matrix $S$ every 1,000 iterations up to 10,000 iterations. We set the kernel size as 0.05 and the loss weight as $\lambda=0.2$.
@@ 4.2 Result Analysis
Comparison on the Mip-NeRF 360 Dataset In our experiments, we trained models on data downsampled by a factor of eight and rendered images at multiple resolutions (1/8, 1/4, 1/2, and full). As shown in the Tab. 1, our method outperforms the baseline models in rendering images at both high and low resolutions. Fig. 4 shows our approach yields high-fidelity images without high-frequency artifacts, contrasting with Mip-NeRF 360 [2] and Zip-NeRF [3], which struggle at higher resolutions due to MLP limitations. Furthermore, traditional 3DGS [9] often suffers from degradation artifacts, and 3DGS + EWA [32] methods exhibit high-frequency artifacts. Our approach avoids these issues, more accurately represents the ground truth, and reduces blurred artifacts seen in Mip-Splatting [30].
Comparison on the Villa Dataset In the Villa Dataset experiment, we compared our model’s training progression against Mip-Splatting [30], with both models trained at the original resolution. Results in Fig. 5 show the performance of both models across 100, 900, 2k, and 5k iterations. Our model demonstrated significant improvements by the 900th iteration, whereas scenes from Mip-Splatting [30] remained blurry and of lower quality. Even after 5k iterations, Mip-Splatting failed to achieve the detail level of our model at 900 iterations, especially in fine features such as roofs, windows, and walls.
Evaluation on the Large Scene Dataset In our study, we addressed the challenges of reconstructing large scenes like small towns or city-scale environments, which are unmanageable for traditional 3DGS-based [9] and NeRF-based [15] models due to memory constraints and long optimization times. We used the Mill-19 Rubble scene [23], which had excessively noisy point clouds requiring reprocessing and selective image filtering. We subdivided the scene, which contained over 1,700 images, into 64 blocks. Each block was independently trained. This reduced memory demands and allowed efficient parallel training in just 20 minutes. As shown in Fig. 6, our reconstruction results show the seamless reassembly of all blocks, which preserves the continuity of the large-scale scene. This method contrasts with previous models, which failed to directly reconstruct large scenes and compromised on reconstruction quality by randomly selecting a subset of images for training.
@@
Figure 7: Ablation of Gaussian Sphere Constraints Strategy. We present an ablation study of our model trained on the Villa scene, comparing results with and without the application of the Gaussian Sphere Constraints Strategy. We present different viewing angles of the training Gaussians and the 3D scene. This strategy reduces the “air walls” problem.
@@ 4.3.1 Patch Attention Enhancement
Omitting the $L_{Frequency}$ and $L_{Edge}$ from the Patch Attention strategy leads to a significant reduction in rendering quality. To quantitatively evaluate this impact, we produced Table. 2, which compares performance metrics with and without this enhancement. The results clearly show that applying the Patch Attention strategy leads to consistent improvements across all metrics, significantly enhancing the model’s ability to generate sharper and more detailed renderings by focusing on edge information.
@@ 4.3.2 Gaussian Sphere Constraints
We assessed the importance of Gaussian Sphere Constraints by removing them from our model. As shown in Fig. 7, models rendered without these constraints exhibit oversized Gaussian spheres, which result in information loss and reduce the overall quality of the renderings. In 3D scenes, these oversized spheres often create “air walls” in detail-heavy areas. Implementing Gaussian Sphere Constraints allows us to effectively control the growth and size of these spheres, enhancing detailed depiction within the scene. Fig. 7 clearly demonstrates the loss of detail in models rendered without this strategy. These images highlight how the constrained Gaussian spheres maintain finer details, leading to more precise and realistic renderings. Additionally, as indicated in Table. 2, the inclusion of Gaussian Sphere Constraints significantly improves performance metrics.
@@ 4.3.3 Subdivision-Based Reconstruction of Large Scenes
In this work, we are unable to conduct an ablation study for our “Subdivision-Based Reconstruction of Large Scenes” due to the inability of existing 3DGS-based methods [9, 30, 14, 5, 8, 7] to handle large-scale scenes without out-of-memory errors. Our subdivision strategy is essential for enabling 3DGS training on such scenes—a task infeasible for current methods without significant modifications. As this strategy is not a mere enhancement but a prerequisite for large-scale training, removing it for ablation would render the application of 3DGS on large scenes impossible.
@@ 4.4 Limitation
Our Patch Attention Enhancement increases memory demands and potentially causes out-of-memory errors during training, especially in complex scenes. Future model iterations could explore more efficient computational strategies to alleviate this issue. Additionally, assembling large scenes reveals overlaps at block boundaries, causing visual discontinuities. A future algorithm could remove Gaussian spheres at these boundaries to improve scene continuity.
@@ 5 Conclusion
In this paper, we present GaussianFocus, an enhanced model built upon traditional 3D Gaussian Splatting. It features three key innovations: Patch Attention Enhancement, Gaussian Constraints Strategy, and the subdivision of large-scale scenes into manageable blocks for individual training. These innovations aim to refine detail representation, enhance reconstruction quality and reduce the “air walls” problem. The approach of subdividing large scenes into manageable blocks overcomes the limitations inherent in traditional 3DGS-based methods, which struggle with extensive scenes. Experimental results demonstrate that GaussianFocus competes well with state-of-the-art methods at a single scale and excels across multiple scales, providing superior detail accuracy and reconstruction quality.
{credits}
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
6 Acknowledgement
This preprint has not undergone peer review (when applicable) or any post-submission improvements or corrections. The Version of Record of this contribution is published in Communications in Computer and Information Science (CCIS, volume 2756), and is available online at https://doi.org/10.1007/978-981-95-4097-6_2.